OpenStudy (anonymous):

Simplify: i^67 1) i 2) -1 3) -i 4) 1

OpenStudy (anonymous):

i^4n =1

OpenStudy (anonymous):

So i^4*16 = i^64 = 1

OpenStudy (anonymous):

1*i^2 = 1*-1 =-1

OpenStudy (anonymous):

i^64*i^2 = -1 *i =-i

OpenStudy (anonymous):

so its #3?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So if it was i^72...the answer would still be -i?

OpenStudy (anonymous):

i^4*18 = i^72 = 1 If your power divides by 4 then it's 1. So you get as near as u can. Say it was 74, then u have 1 (the 72) * i^2 (the other 2) =1*-1 = -1.